OpenStudy (anonymous):

A silver piece of art in the shape of a regular heptagon has an apothem length of 27 mm and a side length of 26 mm. What is its area? ____ mm2

OpenStudy (anonymous):

@iGreen.

Directrix (directrix):

The area of a regular polygon is 1/2 * apothem * perimeter. A heptagon has 7 sides making the perimeter7 * 26 Area = (1/2) * 27 * 7 * 26 = ? Crank that out and you'll have the area.

Directrix (directrix):

Question?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

2457?

Directrix (directrix):

Let me check. Just a sec.

Directrix (directrix):

2457 square mm, same as yours